OpenStudy (anonymous):

what is the intergration of 1.

OpenStudy (abb0t):

\(\int dx = x+C\)

OpenStudy (anonymous):

why i dont understand

OpenStudy (abb0t):

integration is the anti-deriviative. If you take the derivative of "x", hence \(\large \frac{d}{dx}(x) = 1\) going backwards (anti-derivative) or integrate, \(\int 1 dx = x\)

OpenStudy (anonymous):

thanks you so much sir